12123333333333333333:

NEED THE ANSWER FAST!! Will give medal. A college is currently accepting students that are both in-state and out-of-state. They plan to accept three times as many in-state students as out-of-state, and they only have space to accept 100 out-of-state students. Let x = the number of out-of-state students and y = the number in-state students. Write the constraints to represent the incoming students at the college. 0 < x ≤ 100 and 0 < y ≤ 300 x > 0 and y > 0 0 < x ≤ 100 and y > 300 0 < x and y < 100

12123333333333333333:

Braden bought a piece of commercial real estate for $101,234. The value of the real estate appreciated at a constant rate per year. The table shows the value of the real estate after the first and second years: Year 1 2 Value (in dollars) $104,271.02 $107,399.15 Which function best represents the value of the real estate after t years? f(t) = 101,234(1.03)t f(t) = 101,234(0.03)t f(t) = 104,271.02(1.03)t f(t) = 104,271.02(0.03)t

12123333333333333333:

Thanks!

Enderdude:

welcome